THE Mg-NC STRETCHING VIBRONIC BANDS OF THE MgNC Ã ^{2}Π – X̃ ^{2}Σ^{+} TRANSITION

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and measured the laser induced fluorescence excitation spectra of the Mg-NC stretching vibronic bands of the Ã ^{2}Π – X̃ ^{2}Σ^{+} transition. In addition to the bands already reported by Wright and Miller^{a}, the 3_{0}^{2} and 3_{0}^{3} bands have been observed for the first time, and the molecular constants of the vibronic levels, (0,0,2) and (0,0,3), in the Ã ^{2}Π state have been newly determined through rotational analysis of the observed bands. Comparing all of the spin-orbit constants, A_{SO}, on the vibrational levels for the two stretching modes, Mg-NC and C-N, obtained in this work and reported previously^{b}, we find opposite dependence of A_{SO} upon the two stretching vibrational modes, i.e. A_{SO} increases with the excitation of the Mg-NC stretching vibrational mode, while it decreases with the excitation of the C-N mode. The experimental finding is supported by an attempt to calculate the dependence of A_{SO} upon the distance along the vibrational modes.

aR. R. Wright and T. A. Miller, J. Mol. Spectrosc. \underline{194}, 219 (1999).

bM. Fukushima and T. Ishiwata, J. Mol. Spectrosc. \underline{216}, 159 (2002).